A  brass  plug has  a  diameter  of 10,000 cm  at  20°C.  To what  temperature  must  it  be  cooled in order to fit  into  a  hole  A  brass  plug has  a  diameter  of 10,000 cm  at  20°C.  To what  temperature  must  it  be  cooled in order to fit  into  a  hole  with a  diameter  of 9,997 cm?  (α  =19×  10–  6)





1
Expert's answer
2021-06-24T11:58:10-0400

By the definition of the linear thermal expansion we have:



"\Delta L = \alpha L_0 \Delta T,"

here, "\Delta L = L - L_0" is the increace in diameter of the brass plug when it is heated, "L_0 = 10 \ m" is the initial diameter of the brass plug at some temperature "T_0""L = 10 \ m" is the diameter of the brass plug at "T = 20 \ ^{\circ}C""\alpha = 19 \cdot 10^{-6} \ ^{\circ}C^{-1}" is the linear expansion coefficient for steel, "\Delta T = T - T_0" is the change in temperature.

From this formula we can find the change in temperature:



"\Delta T = \dfrac{L - L_0}{\alpha L_0},""\Delta T = \dfrac{10 \ m - 9.997 \ m}{19 \cdot 10^{-6} \ ^{\circ}C^{-1} \cdot 9.997\ m} = 15.79 \ ^{\circ}C."

Finally, we can calculate the temperature at which the brass plug will fit exactly into a hole of constant diameter:



"T_0 = T - \Delta T = 20 \ ^{\circ}C - 15.79 \ ^{\circ}C = 4.21 \ ^{\circ}C."

Answer:

"T_0 = 4.21 \ ^{\circ}C."